OpenStudy (anonymous):

A man on a platform notices that a train going in one direction takes 10 seconds to pass him & a train of same length going in the other direction takes 15 seconds to pass him. What is the time taken by the two trains to pass one another if the length of the trains is 200 m each?

OpenStudy (anonymous):

200/10=20 m/s 200/15=40/3 m/s

OpenStudy (anonymous):

USe relative velocity concept

OpenStudy (anonymous):

idk that concept

OpenStudy (anonymous):

time=distance/speed

OpenStudy (anonymous):

time =400/(20+40/3)

OpenStudy (anonymous):

12 sec